There is a number in the box [ ] such that [ ], 24, 9, 12 are in proportion. The number in the box is _____.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

There is a number in the box [ ] such that [ ], 24, 9, 12 are in proportion. The number in the box is 18.

Let us assume the missingnumber be ‘P’,Then, P, 24, 9, 12
P: 24 = 9: 12
P/24 = 9/12

9/12 is further simplified by dividing both numerator and denominator by 3. So, P/24 = 3/4
P = (3 × 24)/4
P = 72/4

P = 18
Therefore, the missing number is 18.

Thanks